OpenStudy (anonymous):

Part 1. Create two radical equations: one that has an extraneous solution, and one that does not have an extraneous solution. Use the equation below as a model. a√x+b+c=d Use a constant in place of each variable a, b, c, and d. You can use positive and negative constants in your equation.

OpenStudy (anonymous):

I need help answering the question

OpenStudy (anonymous):

please someone help have to finish flvs

OpenStudy (jdoe0001):

well... the extraneous would be the one... may take more the non-extraneous is just any equation more or less the extraneous could be \(\bf 3\sqrt{x}=-5\implies \sqrt{x}=-\cfrac{5}{3}\implies x=\left( \cfrac{-5}{3} \right)^2\implies x=\cfrac{(-5)^2}{3^2} \\ \quad \\ x=\cfrac{25}{9}\) now.. why is that extraneous? or strange or odd ? well... because positive square root of a number, will never yield a negative root yet the procedure at \(\sqrt{x}=-\cfrac{5}{3}\) assumes that

OpenStudy (jdoe0001):

as far as a non-extraneous, well, just grab a couple of values get an equation, any will do

OpenStudy (anonymous):

thank you I had no idea even where to start until you helped

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

thx